Write a variable expression that represents each phrase. If necessary, choose a variable to represent the unknown number. (these are 2 different problems)

1. The quotient of a number and 4 more than another number results in 45



2. 14 more than 6 times a number cubed results in 37


Thank you for your help,

Answer by checkley77(12844)

x(y+4)=45
-------------------------------
6x^3+14=37
6x^3=37-14
6x^3=23
x^3=23/6
x=cubert23/6
x=1.565 answer.
Proof:
6*1.56563+14=37
23+14=37
37=37